OpenStudy (anonymous):

Help please Mary has $4.00 in nickels, dimes, and quarters. If she has twice as many dimes as quarters and forty-seven more nickels than dimes, how many coins of each type does she have?

OpenStudy (anonymous):

I just need the answer please

OpenStudy (anonymous):

Solve the following for n, d and q, the number of nickles, dimes and quarters.\[\left\{5n+10d+25q=400,\frac{d}{q}=2,d=n-47\right\}\]\[\{n=53,d=6,q=3\} \]